# K-Function: Joint Pronunciation Transcription and Feedback for Evaluating Kids Language Function

### v2

@@ 1 Introduction
1 in 14 children in the U.S. has a speech or language disorder, such as dyslexia or Autism, that can hinder reading, vocabulary growth, and communication [1, 2, 3]. These delays can affect quality of life, academic success, and future opportunities. Thus, developing a reliable and automatic framework for evaluating kids language function, covering verbal skills, developmental milestones, reading, and comprehension, is essential. It enables early detection of delays and supports educators and specialists in early identification and intervention.
The bottleneck in developing such a framework lies in the accurate transcription of kids speech, which is significantly more challenging than that of adults. This difficulty arises from several unique characteristics of child vocalizations, including higher fundamental frequency, longer phone durations, greater articulatory variability, and persistent data sparsity [4, 5, 6, 7, 8].
Although recent work has produced increasingly robust child-oriented Automatic Speech Recognition (ASR) models [9, 10, 11, 12, 13, 14, 15, 16], most developmentally significant pronunciation errors arise at the sub-word level and are consequently hidden by word-level metrics such as Word Error Rate (WER). Consider the prompt bird /b@rd/: a child who utters /bed/ (“bed”) exhibits a phoneme substitution, while /pb@rd/ (“p-bird”) contains an intrusive plosive. State-of-the-art end-to-end ASR systems often output the canonical word bird for both productions, masking the underlying articulatory deviation. A dedicated sub-word recognizer that reliably recovers phoneme sequences is therefore indispensable for fine-grained kids language function evaluation, where phoneme recognition is the most widely used approach.
Traditional work has primarily focused on fine-tuning children's speech data on pretrained self-supervised speech learning (SSL) models [17, 18, 19, 20, 21, 22, 23]. However, these approaches exhibit poor generalizability to open-domain children's speech. Moreover, directly adopting adult-oriented ASR or phoneme recognition pipelines fails to address allophonic variations and disrupted phoneme-level language structures.
@@
In this work, we present K-Function, a comprehensive framework for evaluating children’s language function. It consists of our newly proposed Kids-Weighted Finite State Transducer (K-WFST), a lightweight and interpretable kids phoneme transcription model—an LLM-boosted automatic language function scoring system, and an feedback component. Together, these modules form a unified assessment framework for generalized language function evaluation in children.
@@ 2.2 K-WFST: Phonetic Recognizer for Disorder Screening
The process for generating this augmented graph is formally detailed in Algorithm 1. The algorithm's inputs are the reference phoneme ID sequence $R$, a hyperparameter $\beta$ that controls the penalty for errors, and SimMatrix [34]. The SimMatrix is a pre-calculated $N\times N$ matrix where $N$ is the number of phonemes in the vocabulary; each entry SimMatrix[$p_{i},p_{s}$] contains a score between 0 and 1 representing the phonetic similarity between phoneme $p_{i}$ and $p_{s}$. The algorithm initializes an empty set of arcs, $L$, and iterates through all start states $i$ and end states $j$ in the reference sequence, skipping the case where $i=j$ to prevent self-loops. From $\beta$, it derives weights for correct transitions ($\alpha$) and errors ($w_{\text{err}}$). It then populates $L$ with arcs for correct paths, substitutions, deletions, and repetitions, with weights determined by $\alpha$, $w_{\text{err}}$, and the similarity scores.
@@ 16           $w_{\text{sub}}\leftarrow w_{\text{err}}\times(1-\text{SimMatrix}[p_{i},p_{s}])$
Furthermore, to control the flexibility of this substitution mechanism, we employ an Adaptive K-Selection strategy. This allows the decoder to operate in two distinct modes depending on the input speech characteristics and the base acoustic model's performance:
@@
When $K=1$, the model is constrained to only consider the most similar phoneme, which is the phoneme itself. When $K=2$, the model is allowed to consider the top two most similar phonemes, adding the extra substitution paths to enhance robustness in more challenging scenarios. This adaptive approach ensures optimal performance across different conditions without sacrificing efficiency.
@@ 3.1.2 Evaluation on Multitudes
As shown in the table, the fine-tuned ``Kids-FT Model" consistently outperforms the ``Base Model" across all reading passages, reaffirming the importance of adaptation to child speech. Crucially, on this more challenging disfluent dataset, the flexible WFST (K=2) configuration consistently yields the lowest Phoneme Error Rate (PER) for the fine-tuned model on every single passage. This result strongly supports our Adaptive K-Selection strategy, demonstrating that allowing greater flexibility for phonetically plausible substitutions is essential for enhancing model robustness and accuracy in complex downstream scenarios.
@@ 4 Conclusion and future work
K-Function unifies robust, child-oriented phoneme recognition via our K-WFST framework with sophisticated LLM reasoning into an end-to-end pipeline. This system converts children's speech into objective scores and valuable insights for language assessment. Its modular architecture shows strong promise for broad deployment across educational and interventional settings. Future work will focus on expanding the framework to multilingual contexts, refining the analysis to finer linguistic units such as syllables, and verifying its long-term impact and fairness through large-scale field studies.

### v3

@@ 1 Introduction
1 in 14 children in the U.S. has a speech or language disorder, such as dyslexia or Autism, that can hinder reading, vocabulary growth, and communication [1, 2, 3, 4]. These delays can affect quality of life, academic success, and future opportunities. Thus, developing a reliable and automatic framework for evaluating kids language function, covering verbal skills, developmental milestones, reading, and comprehension, is essential. It enables early detection of delays and supports educators and specialists in early identification and intervention.
The bottleneck in developing such a framework lies in the accurate transcription of kids speech, which is significantly more challenging than that of adults. This difficulty arises from several unique characteristics of child vocalizations, including higher fundamental frequency, longer phone durations, greater articulatory variability, and persistent data sparsity [5, 6, 7, 8, 9, 10].
Although recent work has produced increasingly robust child-oriented Automatic Speech Recognition (ASR) models [11, 12, 13, 14, 15, 16, 17, 18, 19], most developmentally significant pronunciation errors arise at the sub-word level and are consequently hidden by word-level metrics such as Word Error Rate (WER). Consider the prompt bird /b@rd/: a child who utters /bed/ (“bed”) exhibits a phoneme substitution, while /pb@rd/ (“p-bird”) contains an intrusive plosive. State-of-the-art end-to-end ASR systems often output the canonical word bird for both productions, masking the underlying articulatory deviation. A dedicated sub-word recognizer that reliably recovers phoneme sequences is therefore indispensable for fine-grained kids language function evaluation, where phoneme recognition is the most widely used approach.
Traditional work has primarily focused on fine-tuning children's speech data on pretrained self-supervised speech learning (SSL) models [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. However, these approaches exhibit poor generalizability to open-domain children's speech. Moreover, directly adopting adult-oriented ASR or phoneme recognition pipelines fails to address allophonic variations and disrupted phoneme-level language structures.
@@
In this work, we present K-Function, a comprehensive framework for evaluating children’s language function. It consists of our newly proposed Kids-Weighted Finite State Transducer (K-WFST), a lightweight and interpretable kids phoneme transcription model—an LLM-boosted automatic language function scoring system, and a feedback component. Together, these modules form a unified assessment framework for generalized language function evaluation in children.
@@ 2.2 K-WFST: Phonetic Recognizer for Disorder Screening
The process for generating this augmented graph is formally detailed in Algorithm 1. The algorithm's inputs are the reference phoneme ID sequence $R$, a hyperparameter $\beta$ that controls the penalty for errors, and the SimMatrix [41]. The SimMatrix is a pre-calculated $N\times N$ matrix where $N$ is the number of phonemes in the vocabulary. We construct SimMatrix using a heuristic approach based on eight phonological features like vowel height, voicing. The similarity between $p_{i}$ and $p_{s}$ is computed as the normalized weighted sum of their matching features, yielding a score in $[0,1]$ that captures linguistically grounded proximity. The algorithm initializes an empty set of arcs, $L$, and iterates through all start states $i$ and end states $j$ in the reference sequence, skipping cases where $i=j$ to prevent self-loops. From $\beta$, it derives weights for correct transitions ($\alpha$) and errors ($w_{\text{err}}$). It then populates $L$ with arcs for correct paths, substitutions, deletions, and repetitions, with weights determined by $\alpha$, $w_{\text{err}}$, and the similarity scores.
@@ 16           $w_{\text{sub}}\leftarrow w_{\text{err}}\times(1-\text{SimMatrix}[p_{i},p_{s}])$
Furthermore, to control the flexibility of this substitution mechanism, we employ a Task-dependent K-Selection strategy. This allows the decoder to operate in two distinct modes depending on the input speech characteristics and the base acoustic model's performance:
@@
When $K=1$, the model is constrained to only consider the most similar phoneme, which is the phoneme itself. When $K=2$, the model is allowed to consider the top two most similar phonemes, adding the extra substitution paths to enhance robustness in more challenging scenarios. This task-dependent approach ensures optimal performance across different conditions without sacrificing efficiency.
@@ 3.1.2 Evaluation on Multitudes
As shown in the table, the fine-tuned ``Kids-FT Model" consistently outperforms the ``Base Model" across all reading passages, reaffirming the importance of adaptation to child speech. Crucially, on this more challenging disfluent dataset, the flexible WFST (K=2) configuration consistently yields the lowest Phoneme Error Rate (PER) for the fine-tuned model on every single passage. This result strongly supports our Task-dependent K-Selection strategy, demonstrating that allowing greater flexibility for phonetically plausible substitutions is essential for enhancing model robustness and accuracy in complex downstream scenarios.
@@ 4 Conclusion and future work
K-Function unifies robust, child-oriented phoneme recognition via our K-WFST framework with sophisticated LLM reasoning into an end-to-end pipeline. This system converts children's speech into objective scores and valuable insights for language assessment. Its modular architecture shows strong promise for broad deployment across educational and interventional settings. Notably, K-WFST is inherently language-agnostic, as the SimMatrix relies on universal articulatory features adaptable to any language with defined phonetics. Future work will focus on expanding the framework to multilingual contexts, refining the analysis to finer linguistic units such as syllables, and verifying its long-term impact and fairness through large-scale field studies.